What is the molarity of a sodium hydroxide solution if 27.9 mL of this solution reacts exactly with 22.30 mL of 0.253 M sulfuric acid?
1
Expert's answer
2013-07-16T11:07:58-0400
n(H2SO4) = Cm x V
n(H2SO4) = 0.253 M x 0.02230 L = 0.00564 mol
The balanced equation is:
H2SO4 + 2NaOH = Na2SO4 + 2H2O
the ratio between H2SO4 and NaOH is 1 : 2
n(NaOH) = 0.00564/2 = 0.00282 mol
Cm(NaOH) = n(NaOH)/V(NaOH)
Cm(NaOH) = 0.00282 mol/ 0.0250 L = 0.113 M

Answer: 0.113 M